The length of a rectangle is 5 cm less than 3 times its width. If the perimeter of the rectangle is 46 cm, find the dimensions of the rectangle.
Answer by checkley77(12844)

Sorry. I must have bneen tired & misread the problem.
You're right.
L=3W-5
P=2L+2W
46=2(3W-5)+2W
46=6W-10+2W
46=8W-10
8W=46+10
8W=56
W=56/8
W=7 cm. ANS. FOR THE WIDTH.
L=3*7-5=21-5=16 cm. ANS. FOR THE LENGTH.
PROOF:
46=2*16+2*7
46=32+14
46=46